OpenStudy (anonymous):

Each dimension of a rectangular prism is increased by 10%. What percent of the original volume is the volume of the new prism? 00.1 percent 133.1 percent 121 percent 1 percent

OpenStudy (phi):

if you know the ratio of lengths of 2 similar prisms, then their volumes will have a ratio of the lengths cubed: (L1/L2)^3= V1/V2 where L means length and V means volume for 10% bigger length L + 10% of L= L+ L/10 or L+0.1*L or L(1+0.1) or L*1.1 I would find (1.1/L)^3 divided by (1/L)^3 the L^3 parts will cancel, so the fraction is 1.1^3/1^3 or just 1.1^3 change to a percent by multiplying by 100

OpenStudy (anonymous):

b

OpenStudy (phi):

133.1%

OpenStudy (anonymous):

thnks